OpenStudy (anonymous):

the height of a right circular cylinder is 10cm and the diameter of the base is 18 cm find the Surface area

OpenStudy (anonymous):

height (h) = 10 cm radius (r) = 18/2 = 9 cm Curved Surface area = 2*pi*r*h = 2 x 3.14 x 9 x 10 = 565.2 sq cm. Total Surface Area = 2*pi*r*h + 2*pi*r*r = 565.2 + 508.7 = 1073.9 sq cm

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

Cylinder Surface Area = 2 x (3.1416 x Radius2) + (2 x 3.1416 x Radius x Length) where diameter is twice of radius you can do it now :D